# MultiTok: Variable-Length Tokenization for Efficient LLMs Adapted from LZW Compression

### v2

@@ Abstract
Large language models have drastically changed the prospects of AI by introducing technologies for more complex natural language processing. However, current methodologies to train such LLMs require extensive resources including but not limited to large amounts of data, expensive machinery, and lengthy training. To solve this problem, this paper proposes a new tokenization method inspired by universal Lempel-Ziv-Welch data compression that compresses repetitive phrases into multi-word tokens. With MultiTok as a new tokenizing tool, we show that language models are able to be trained notably more efficiently while offering a similar accuracy on more succinct and compressed training data. In fact, our results demonstrate that MultiTok achieves a comparable performance to the BERT and GPT-2 standards as both a stand-alone tokenizer and an add-on to existing tokenizers while also providing close to 2.5x faster training with more than 30% less training data.
Index Terms:
compression, efficiency, language models, universal compression, multi-word tokenization.
@@ I Introduction
Tokenization is an essential component of many current large language pipelines [9, 10, 11], which directly affects the volume of training data and training complexity. Some examples of tokenization algorithms in the literature are variants of the WordPiece [12] and Byte Pair Encoding [13, 14]. These algorithms are building blocks for generating the state-of-the-art tokens utilized in the Bidirectional Encoder Representations from Transformers (BERT) [15] and Generative Pre-trained Transformer 2 (GPT-2) [16] models.
@@
One of the most well-known lossless compression algorithms is the Lempel-Ziv-Welch (LZW) algorithm [24]. The LZW algorithm is a universal compression algorithm that works by compressing repeated patterns of bits within a dataset. As such, it achieves a high compression ratio and performs better with larger amounts of data [25]. In this paper, we introduce MultiTok, a new tokenization method inspired by the LZW compression algorithm for the efficient LLM training tasks. We focus on LLMs where the inputted data is first tokenized utilizing an encoder, extracted into embedding vectors representing the semantic and contextual information, and finally fed into a transformer model for text classification. MultiTok provides a novel variable-length tokenizer in the sense that each token can represent a variable number of words. The advantages of MultiTok are it (i) dynamically compresses the necessary training data by $33\%$ (ii) allows the LLM to be trained close to 2.5x times faster and (iii) maintains performance comparable to the baseline BERT and GPT-2 token standards. Our tokenizing schemes can mark the beginning of the use of information-theoretic approaches to provide efficient, secure, and robust LLM systems. The code and data can be accessed at: https://github.com/noelkelias/multitok.
@@ III MultiTok: A Variable-Length Tokenization
An overview of the MultiTok tokenization scheme is illustrated as Algorithm 1. The input of the algorithm is the training dataset $\mathcal{X}$, an initial dictionary of all possible single-word tokens $D$, and the maximum number of words that a MultiTok token can represent as $w$. In Line 3, we initialize the MultiTok vector $T(x)$ for sample $x$ and the current length of our dictionary as $idx$. Through Lines 4 and 5, we iterate through each word within $x$. For each word, we look $w$-word ahead to form a multi-word phrase of maximum length $w$, starting from the shortest phrase. When we find the smallest multi-word phrase that is not already in our dictionary $\mathcal{D}$, we will add it to the dictionary (lines 6 and 7). Then, we add the current largest known multi-word token to our MultiTok vector $T(x)$ and move on to encode the next available word (Line 7). If we have reached the end of our sample, we add the index of the remaining phrase to our MultiTok vector in Line 10. Finally, we add the vector of tokens representing the sample $x$ to the set $\mathcal{T}$.
We note that $x[i{:}j]$ is the sub-vector of $x$ with the elements from indices $i$ to $j$. Further, the operator +={+=}+ = denotes appending an element to the end of a vector. During inference, for tokenization, we start from the first word of a sample, and identify the largest match in the dictionary with the length at most $w_{\text{test}}$ (called testing window). We note that $w_{\text{test}}$ can be different than $w$ (called training window) in Algorithm 1.
1:$\mathcal{X},D,w$ 2:$\mathcal{T}=\{T(x)\}_{x\in\mathcal{X}}$ 3:$\mathcal{T}=\{\}$ 4:$idx=|\mathcal{D}|+1$ 5:for $x\in\mathcal{X}$ do 6:   $i=1,T(x)=[{}]$ 7:   while $i\leq|x|$ do 8:      for $j\in[i,\text{min}(i+w,|x|)]$ do 9:         if $x[i:j]\notin D$ then 10:           $D(x[i:j])=idx,\;idx=idx+1$ 11:           $T(x){+\!=}D(x[i:j-1]),\;i:=j$ 12:           break           13:         if $j=|x|$ then $T(x){+\!=}D(x[i:j])$                 14:      $\mathcal{T}=\mathcal{T}\cup T(x)$
@@
Lemma 1 describes how much the compression ratio increases after the proposed frequency-based post-processing.
Lemma 1.
@@ IV Simulations Results
In this section, we demonstrate the empirical performance of our tokenizer, MultiTok, with respect to the utility and efficiency goals described by Definitions 1 and 2. Further, we show various goals such as utility, accuracy, and training time can compete, and we demonstrate some parameters that can offer a good trade-off among them.
@@ IV-A Experimental Setup
The general pipeline includes tokenizing the training samples with the MultiTok, BERT, and GPT-2 tokenization schemes. These tokens were then assigned corresponding random vector embedding and inputted to train a selected model for the textual binary classification task.
@@
Evaluation. All our results are the averages over 20 trials with the same parameters. Specifically, we record the model’s average training loss, training accuracy, and testing accuracy per epoch. In addition, the compression ratio as well as the average area under the curve (AUC) given by the receiver operating characteristic curve (ROC) for each test were recorded and graphed. Our results for MultiTok were compared with BERT tokenization [15], GPT-2 [16] tokenization, and single-word tokenization where each unique word is considered as a token. For all these schemes, we use random vectors for embedding of each token. We note that the BERT and GPT-2 tokenizers are well-known off-the-shelve tokenizers utilized for LLM training today, and thus we targeted them as baselines in our experiments.
Our first round of experiments are focused on studying the performance MultiTok algorithm for various sizes of the training window and the testing window. In the second experiment, we apply MultiTok encoding on the BERT and GPT-2 tokens to have a cascade of two tokenization schemes, where MultiTok takes BERT or GPT-2 tokens as units of data for compression. Finally, we apply the frequency-based post processing on the MultiTok tokens for further improvements. We note that the title of rows in the table can be read as the type of tokenization, the percentage of training data to which the tokenizer is applied, and the training and testing window sizes.
@@
Figure 3: MultiTok demonstrates an average training time ($C(\epsilon)|\epsilon=0.01$) that is almost 2-3 times faster than the BERT & GPT-2 standards.
@@ IV-C Combining MultiTok with Existing Tokenizers
In this experiment, we used MultiTok tokenization to further compress the repetitive sequences within BERT and GPT-2 tokens. This consequently reduces the training data size while achieving similar performance to that of normal uncompressed BERT & GPT-2 tokens. We observed that we could compress the training data that is tokenized via BERT by $26\%$ and via GPT-2 by $38\%$ using MultiTok, while achieving the highest accuracies of any downstream model with $75.3\%$ and $74.5\%$ respectively. However, when we further reduced training data size, a higher compression is obtained at the cost of slight degradation in the performance. This suggests a notion of using a ‘golden ratio’ of MultiTok compression to maximize efficiency and performance. This ‘golden ratio’ may vary based on the repetition and succinctness found within different datasets.
Next, we provide an empirical training convergence analysis, which shows that MultiTok not only performs as good as the BERT & GPT-2 standards in terms of accuracy and AUC, but also directly results in a faster training process for the downstream models.
We observe in Fig. 2 that applying MultiTok tokenization to existing BERT and GPT-2 tokens not only reduces the size of the training data, but also allows the model to train faster. Specifically, the combination of MultiTok and BERT or GPT-2 tokenization achieves an average training time (as defined in Definition  2) of epoch 11 rather than epoch 20 (almost $2x$ faster) as shown in Fig. 3. Thus, these experiments suggest that MultiTok can be cascaded with existing tokenizers, like BERT and GPT-2, for smaller training data sizes, more efficient training, and improved downstream model performance.
@@ IV-D MultiTok Tokenization with Post-Processing
Our results in this experiment indicates that the post processing of MultiTok tokenization with a frequency analysis component dramatically increases the performance of the resulting tokens. We remind that without this post processing, there is a significant decrease in the model’s performance when the full capacity of MultiTok is utilized (when MultiTok is applied to the entire dataset with maximum window sizes), see Table I. However, by pruning less-utilized tokens, we were able to circumvent any impact to accuracy, as shown by the last set of experimental results. Specifically for experiment “MultiTok, 100%, 2-2, $\geq$2” (MultiTok is applied to the entire dataset with a training window size of 2), we see that by simply removing all multi-word tokens that only appear once in the training data, we get an accuracy of around $73\%$. While the training data size of the initial MultiTok experiment (MultiTok, 100%, 2-2,) increases by $6\%$, the resulting more succinct tokens provide an accuracy comparable to the BERT & GPT-2 baseline standards while still reducing the original training data size by $34\%$.
According to Fig. 2, using these post-processed MultiTok tokens can result in a comparable accuracy up to $2.5x$ faster than when using BERT or GPT-2 tokens. Most impressively, at epoch 8 in Fig. 2, the loss of this stand-alone post-processed MultiTok experiment has almost achieved its stable minimum value, being at least 0.4 lower than the training loss for the baselines at the same epoch. In addition, as shown in Fig. 3, the average training time for MultiTok based experiments is 8-12 epochs, which is significantly lower than the other counterpart tokenizers. Significantly, MultiTok has the lowest compression ratio out of the selected experiments, where the MultiTok tokens not only compressed the inputted training data by a factor of around $3$, but also helped the model learn the training data almost $2.5x$ faster. This confirms that with MultiTok tokenization, the model is able to learn the training data and existing tokens faster and more efficiently as opposed to the baseline tokenization schemes.
@@ V Conclusion & Future Work
We proposed MultiTok, a novel multi-word tokenizer to notably compress repetitive words or phrases within a training dataset without harming the performance of downstream models. The proposed scheme can be used stand-alone and applied directly on words of text-based data, or it can be followed by an off-the-shelf tokenizer to improve efficiency and performance. Specifically, we demonstrate being able to compress training data by up to $33\%$ and train the downstream model $2.5x$ faster while maintaining comparable performance to the BERT and GPT-2 standards. A promising next step is to extend MultiTok to be used for complex NLP tasks like text generation and even other data types including videos. We hope that this paper marks the start of utilizing such data compression techniques to create more efficient LLMs of the future.

### v3

@@ Abstract
Large language models have drastically changed the prospects of AI by introducing technologies for more complex natural language processing. However, current methodologies to train such LLMs require extensive resources including but not limited to large amounts of data, expensive machinery, and lengthy training. To solve this problem, this paper proposes a new tokenization method inspired by universal Lempel-Ziv-Welch data compression that compresses repetitive phrases into multi-word tokens. With MultiTok as a new tokenizing tool, we show that language models are able to be trained notably more efficiently while offering a similar accuracy on more succinct and compressed training data. In fact, our results demonstrate that MultiTok achieves a comparable performance to the BERT and GPT standards as both a stand-alone tokenizer and an add-on to existing tokenizers while also providing close to 2.5x faster training with more than 30% less training data.
@@ I Introduction
Tokenization is an essential component of many current large language pipelines [20, 17, 10], which directly affects the volume of training data and training complexity. Some examples of tokenization algorithms in the literature are variants of the WordPiece [30] and Byte Pair Encoding [8, 21]. These algorithms are building blocks for generating the state-of-the-art tokens utilized in the Bidirectional Encoder Representations from Transformers (BERT) [5] and Generative Pre-trained Transformer 2 (GPT) [19] models.
@@
One of the most well-known lossless compression algorithms is the Lempel-Ziv-Welch (LZW) algorithm [28]. The LZW algorithm is a universal compression algorithm that works by compressing repeated patterns of bits within a dataset. As such, it achieves a high compression ratio and performs better with larger amounts of data [11]. In this paper, we introduce MultiTok, a new tokenization method inspired by the LZW compression algorithm for the efficient LLM training tasks. We focus on LLMs where the inputted data is first tokenized utilizing an encoder, extracted into embedding vectors representing the semantic and contextual information, and finally fed into a transformer model for text classification. MultiTok provides a novel variable-length tokenizer in the sense that each token can represent a variable number of words. The advantages of MultiTok are it (i) dynamically compresses the necessary training data by $33\%$ (ii) allows the LLM to be trained close to 2.5x times faster and (iii) maintains performance comparable to the baseline BERT and GPT token standards. Our tokenizing schemes can mark the beginning of the use of information-theoretic approaches to provide efficient, secure, and robust LLM systems. The code and data can be accessed at: https://github.com/noelkelias/multitok.
@@ III MultiTok: A Variable-Length Tokenization
An overview of the MultiTok tokenization scheme is illustrated as Algorithm 1. The input of the algorithm is the training dataset $\mathcal{X}$, an initial dictionary of all possible single-word tokens $D$, and the maximum number of words that a MultiTok token can represent as $w$. In Line 3, we initialize the MultiTok vector $T(x)$ for sample $x$ and the current length of our dictionary as idxidx. Through Lines 4 and 5, we iterate through each word within $x$. For each word, we look $w$-word ahead to form a multi-word phrase of maximum length $w$, starting from the shortest phrase. When we find the smallest multi-word phrase that is not already in our dictionary $\mathcal{D}$, we will add it to the dictionary (lines 6 and 7). Then, we add the current largest known multi-word token to our MultiTok vector $T(x)$ and move on to encode the next available word (Line 7). If we have reached the end of our sample, we add the index of the remaining phrase to our MultiTok vector in Line 10. Finally, we add the vector of tokens representing the sample $x$ to the set $\mathcal{T}$.
We note that x[i:j]x[i{:}j] is the sub-vector of $x$ with the elements from indices $i$ to $j$. Further, the operator +={+=} denotes appending an element to the end of a vector. During inference, for tokenization, we start from the first word of a sample, and identify the largest match in the dictionary with the length at most $w_{\text{test}}$ (called testing window). We note that $w_{\text{test}}$ can be different than $w$ (called training window) in Algorithm 1.
@@
0:  $\mathcal{X},D,w$
0:  $\mathcal{T}=\{T(x)\}_{x\in\mathcal{X}}$
1:  $\mathcal{T}=\{\}$
2:  $idx=|\mathcal{D}|+1$
3:  for $x\in\mathcal{X}$ do
4:   i=1,T(x)=[]i=1,T(x)=[{}]
5:   while $i\leq|x|$ do
6:    for $j\in[i,\text{min}(i+w,|x|)]$ do
7:     if $x[i:j]\notin D$ then
8:      D(x[i:j])=idx,idx=idx+1D(x[i:j])=idx,\;idx=idx+1
9:      T(x)+=D(x[i:j1]),i:=jT(x){+\!=}D(x[i:j-1]),\;i:=j
10:      break
11:     end if
12:     if $j=|x|$ then
13:      $T(x)\mathrel{+}=D(x[i:j])$
14:     end if
15:    end for
16:    $\mathcal{T}=\mathcal{T}\cup T(x)$
17:   end while
18:  end for
@@
Lemma III.1 describes how much the compression ratio increases after the proposed frequency-based post-processing.
Lemma III.1.
@@ IV Simulations Results
In this section, we demonstrate the empirical performance of our tokenizer, MultiTok, with respect to the utility and efficiency goals described by Definitions II.1 and II.2. Further, we show various goals such as utility, accuracy, and training time can compete, and we demonstrate some parameters that can offer a good trade-off among them.
@@ IV-A Experimental Setup
The general pipeline includes tokenizing the training samples with the MultiTok, BERT, and GPT tokenization schemes. These tokens were then assigned corresponding random vector embedding and inputted to train a selected model for the textual binary classification task.
@@
Our goal with the simulation was to ensure we most directly evaluate the effect of different tokenizations on a model’s performance. The best way to do this is to eliminate possible confounding variables outside tokenization that could affect a model’s performance. Utilizing more complex architectures in smaller scale experiments could overcompensate a model’s performance for bad tokenization [32]. In addition, the performance of more complex classification tasks are often reflective of the underlying architecture rather than the inputted tokens [23, 20]. The alternative was to train a LLM with vast amounts of training data and billions of parameters, which would negate immediate architectural and other confounding advantages, thereby focusing more directly on inputted training data (tokenizers). For this reason, we used a simple transformer model and tested on a basic classification dataset.
Evaluation. All our results are the averages over 20 trials with the same parameters. Specifically, we record the model’s average training loss, training accuracy, and testing accuracy per epoch. In addition, the compression ratio as well as the average area under the curve (AUC) given by the receiver operating characteristic curve (ROC) for each test were recorded and graphed. Our results for MultiTok were compared with BERT tokenization [5], GPT [19] tokenization, and single-word tokenization where each unique word is considered as a token. For all these schemes, we use random vectors for embedding of each token. We note that the BERT and GPT tokenizers are well-known off-the-shelve tokenizers utilized for LLM training today, and thus we targeted them as baselines in our experiments. Thus, for simulation results, if the proposed MultiTok tokenization strategy was able to perform better than these standards, it would be able to perform better than most other tokenization schemes. For this reason, an extensive comparison of more tokenization schemes, while interesting, would only provide marginal benefits.
Our first round of experiments are focused on studying the performance MultiTok algorithm for various sizes of the training window and the testing window. In the second experiment, we apply MultiTok encoding on the BERT and GPT tokens to have a cascade of two tokenization schemes, where MultiTok takes BERT or GPT tokens as units of data for compression. Finally, we apply the frequency-based post processing on the MultiTok tokens for further improvements. We note that the title of rows in the table can be read as the type of tokenization, the percentage of training data to which the tokenizer is applied, and the training and testing window sizes.
@@
Figure 3: MultiTok demonstrates an average training time ($C(\epsilon)|\epsilon=0.01$) that is almost 2-3 times faster than the BERT & GPT standards.
@@ IV-C Combining MultiTok with Existing Tokenizers
In this experiment, we used MultiTok tokenization to further compress the repetitive sequences within BERT and GPT tokens. This consequently reduces the training data size while achieving similar performance to that of normal uncompressed BERT & GPT tokens. We observed that we could compress the training data that is tokenized via BERT by $26\%$ and via GPT by $38\%$ using MultiTok, while achieving the highest accuracies of any downstream model with $75.3\%$ and $74.5\%$ respectively. However, when we further reduced training data size, a higher compression is obtained at the cost of slight degradation in the performance. This suggests a notion of using a ‘golden ratio’ of MultiTok compression to maximize efficiency and performance. This ‘golden ratio’ may vary based on the repetition and succinctness found within different datasets.
Next, we provide an empirical training convergence analysis, which shows that MultiTok not only performs as good as the BERT & GPT standards in terms of accuracy and AUC, but also directly results in a faster training process for the downstream models.
We observe in Fig. 2 that applying MultiTok tokenization to existing BERT and GPT tokens not only reduces the size of the training data, but also allows the model to train faster. Specifically, the combination of MultiTok and BERT or GPT tokenization achieves an average training time (as defined in Definition  II.2) of epoch 11 rather than epoch 20 (almost $2x$ faster) as shown in Fig. 3. Thus, these experiments suggest that MultiTok can be cascaded with existing tokenizers, like BERT and GPT, for smaller training data sizes, more efficient training, and improved downstream model performance.
@@ IV-D MultiTok Tokenization with Post-Processing
Our results in this experiment indicates that the post processing of MultiTok tokenization with a frequency analysis component dramatically increases the performance of the resulting tokens. We remind that without this post processing, there is a significant decrease in the model’s performance when the full capacity of MultiTok is utilized (when MultiTok is applied to the entire dataset with maximum window sizes), see Table I. However, by pruning less-utilized tokens, we were able to circumvent any impact to accuracy, as shown by the last set of experimental results. Specifically for experiment “MultiTok, 100%, 2-2, $\geq$2” (MultiTok is applied to the entire dataset with a training window size of 2), we see that by simply removing all multi-word tokens that only appear once in the training data, we get an accuracy of around $73\%$. While the training data size of the initial MultiTok experiment (MultiTok, 100%, 2-2,) increases by $6\%$, the resulting more succinct tokens provide an accuracy comparable to the BERT & GPT baseline standards while still reducing the original training data size by $34\%$.
According to Fig. 2, using these post-processed MultiTok tokens can result in a comparable accuracy up to $2.5x$ faster than when using BERT or GPT tokens. Most impressively, at epoch 8 in Fig. 2, the loss of this stand-alone post-processed MultiTok experiment has almost achieved its stable minimum value, being at least 0.4 lower than the training loss for the baselines at the same epoch. In addition, as shown in Fig. 3, the average training time for MultiTok based experiments is 8-12 epochs, which is significantly lower than the other counterpart tokenizers. Significantly, MultiTok has the lowest compression ratio out of the selected experiments, where the MultiTok tokens not only compressed the inputted training data by a factor of around $3$, but also helped the model learn the training data almost $2.5x$ faster. This confirms that with MultiTok tokenization, the model is able to learn the training data and existing tokens faster and more efficiently as opposed to the baseline tokenization schemes.
@@ V Conclusion & Future Work
We proposed MultiTok, a novel multi-word tokenizer which compresses repetitive words and phrases within a training dataset without harming the performance of downstream models. The proposed scheme can be used stand-alone and applied directly on words of text-based data, or it can be followed by an off-the-shelf tokenizer to improve efficiency and performance. Specifically, we demonstrate being able to compress training data by up to $33\%$ and train the downstream model $2.5x$ faster while maintaining comparable performance to the BERT and GPT standards. A promising next step is to extend MultiTok to be used for complex NLP tasks like text generation and even other data types like video. We hope that this paper marks the start of utilizing such data compression techniques to create more efficient LLMs of the future.